Use the following information to answer the next 3 questions: Caroline Garcia, a market researcher, is looking at the combination of products consumers purchase in the same shopping trip. Specifically, she is interested in looking at what goods consumers also purchase when they purchase cigarettes, particularly whether they also purchase milk at that time. From previous research, she knows that 47% of a particular store’s shoppers purchase milk. She expects consumers who purchase cigarettes to have the same probability of purchasing milk. She watches shoppers purchasing cigarettes (smokers) for one day and records 72 who purchase milk and 128 who do not.

11. What is the test statistic for this hypothesis test?

a. -3.14

b. -3.34

c. +3.14

d. +3.34
Answer #1

= x / n = 72 / 200 = 0.36

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.36 - 0.47 / [(0.47 * 0.53) / 200]

= -3.14